# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ 5 Conclusion
This work presents the first systematic investigation of scaling laws for Ephys foundation models (EFMs), introducing DIVER-1, a family of EEG/iEEG foundation models ranging from 13M to 1.82B parameters. Our analysis reveals that EFMs follow data-constrained scaling laws with critical domain-specific characteristics that fundamentally differ from language domains.
Through unprecedented scaling across data volume, model capacity, and computational resources, DIVER-1 achieves state-of-the-art performance on various iEEG and EEG tasks. Our novel architectural innovations including any-variate attention mechanisms, sliding temporal conditional positional encoding, and multi-domain reconstruction heads further enhance model performance.
DIVER also exhibits strong generality. It offers patch-size variants to accommodate different temporal scales across downstream tasks. Importantly, while it can exploit spatial location features, it is designed to function robustly even in the absence of some channel position. Unlike PoPT, BrainBERT, and Brant, which are trained exclusively on SEEG, DIVER is pretrained on both ECoG and SEEG while distinguish them by subtype embedding, and can therefore be applied to both. This versatility makes it a robust foundation model for a wide range of downstream applications.
Also, to better leverage the potential of large EFMs, future work should develop more sophisticated finetuning methodologies, including cross-subject learning frameworks that can handle heterogeneous electrode configurations across subjects, and data-efficient approaches such as LoRA that can effectively adapt large models without overfitting to limited subject-specific data.
Looking forward, these findings establish the foundation for a new generation of neuroscience AI applications by demonstrating that Ephys requires tailored scaling strategies rather than direct application of scaling laws from other domains. While the specific scaling law parameters and optimal ratios we derived are tied to our particular architectural choices and pretext task, the fundamental insight that data-constrained scaling necessitates prioritizing training duration over model size represents a broadly applicable principle for EFM development.
@@ A.1 Ephys Data Foundation Models
Ephys decoding has progressed from pipelines that coupled hand-crafted features with classical classifiers to end-to-end deep architectures that learn task-relevant representations directly from raw signals. Early EEG studies relied on feature engineering (e.g., band-power, CSP) paired with SVMs or LDA; subsequent work introduced convolutional backbones and sequence models that absorb spectral–temporal patterns with minimal preprocessing. Building on advances in self-supervised learning (SSL) and “foundation” paradigms, recent efforts pretrain large models on heterogeneous, weakly labeled or unlabeled EEG/iEEG corpora and adapt them to diverse downstream tasks (e.g., event detection, cognitive state decoding, BCI control). Representative lines include large-scale EEG transformers and montage-aware encoders (e.g., EEGFormer(chen2024eegformer), NEURO-GPT(cui2024neurogpt), LaBraM(jiang2024labram)), intracranial representation learners (e.g., BrainBERT(wang2023brainbert), foundation models for intracranial neural signals), and broader neuro-sequence backbones emphasizing transfer and robustness (e.g., CBraMod(wang2024cbramod)).
@@ A.2 Neural Scaling Law
Kaplan (kaplan2020scaling) scaling law provides a principled framework for predicting model performance as a function of model size and dataset size for large language models. They demonstrated that language model cross-entropy loss follows smooth power-law relationships with respect to model parameters ($N$) and training data ($D$). Concretely, they proposed a relation of the form
@@
Data-constrained scaling law. In scientific domains, including Ephys, unique high-quality data are inherently limited. To address this, data_constrained_scaling_muennighoff2023scaling extended the scaling framework to data-constrained regimes, where models must repeatedly train on the same corpus for multiple epochs. They proposed a modified law
@@
Compared to the Chinchilla law (hoffmann2022chinchilla), which assumes abundant data and one training epoch, this formulation makes epoch count itself a central axis of scaling. This perspective is crucial for domains constrained by limited samples, such as EEG and iEEG, and it guides how compute should be allocated between model size and additional training passes.
Finally, the data-constrained scaling law also connects scaling analysis to more fundamental theory. Recent work has suggested that the exponents governing power-law scaling, such as $\alpha$ and $\beta$, are intimately connected to the intrinsic dimension of the underlying data manifold (Sharma2022Scaling). In particular, it has been argued that the data exponent $\beta$ can be interpreted in terms of an effective dimension $d$ via an approximate inverse relation of the form:
@@ B.2 Pretraining Dataset
AJILE12 (peterson2022ajile12) ECoG 12 124,423 1,282 1,000
@@
TUEG (obeid2016tueg) EEG 10,874 422,036 23,178 250–512
HBN (shirazi2024hbn) EEG 2,782 61,703 572 500
NCHSDB (lee2022nsrr) EEG 3,673 163,146 26,055 256–512
PEERS (peers) EEG 364 870,447 6,964 500
@@ B.3 Downstream Task and Dataset Overview
iEEG tasks. We evaluated on 15 tasks from the Neuroprobe (LITE) benchmark (2024neuroprobe), including visual perception (frame brightness, optical flow, face detection), auditory processing (volume, pitch, delta volume), and language processing (speech decoding, word prediction, onset detection, part-of-speech tagging). The Neuroprobe dataset contains depth electrode recordings from 6 subjects with 109-120 channels per subject, originally sampled at 2048Hz and was resampled to 500Hz to match our pretraining configuration. There are both binary-label and multi-label task options for Neuroprobe.  [^3] The multi-label setting is currently available only in the released code and has not yet been documented in the paper. In the multi-label configuration, the speech, onset, and head word position tasks remain binary, the part-of-speech task uses 6 labels, and the remaining tasks use 3 labels. Throughout this paper, unless explicitly stated otherwise as multi-label, all reported Neuroprobe results correspond to the binary-label setting. Also, we evalauted on the seizure detection test on MAYO dataset (modified the dataset in kaggle challenge (bbrinkm2014upenn)). The original dataset consists of 1 s samples, but to match the minimum patch length of Brant, one of the baseline models, we concatenated samples in temporal order to create 6 s samples. We evaluated the model separately for each of the 8 participants. To address the issue of having more test data than training data, we swapped the train and test sets for each participant.
EEG tasks. We evaluated on three EEG benchmarks: FACED (emotion recognition from 32-channel EEG, 9-class classification, 123 subjects)(chen2023faced), PhysioNet-MI (motor imagery with 64 channels, 4-class classification, 109 subjects)(goldberger2000physionet), and MentalArithmetic (mental stress detection with 20 channels reduced to 19, 2-class classification, 36 subjects)(zyma2019electroencephalograms). Sampling rates were standardized to 500Hz across all datasets to ensure consistency with our pretraining setup.
@@ B.4 Pretraining Setup and Model Scaling
We varied the model size by modifying the hidden dimension of the transformer, resulting in sizes of 13M, 51M, 115M, 203M, 813M, 1.83B parameters, while keeping the depth fixed at 12 layers. This capacity adjustment leverages the benefits of $\mu$ parameterization for stable training across different model sizes. DIVER-1 was implemented on the Python 3.12.3 and Pytorch 2.6.0 + cuda version 12.4. To enhance training efficiency, we employed DeepSpeed ZeRO Stage 2, BF16 precision. Optimization was performed using a custom implementation of the DeepSpeed’s MuAdam optimizer (muparamyang2022tensor) with utilizing DeepSpeed’s FusedAdam backend (rasley2020deepspeed) for computational efficiency and learning rate calibration. A cosine annealing learning rate scheduler with warm-up restarts was applied, with cycle length matching the total training steps and minimum learning rate set to 0.01× the initial rate.
@@ B.5 DIVER Architecture and Pretext Task Hyperparamter Setting
Positional & Embedding Channel type dimension $d_{\text{model}}/4$
@@
Stage2: Optuna Optimization We further refined the hyperparameters using Optuna(optuna) or bayesian hyperparameter optimization. The search space was defined as $\pm 1$ order of magnitude around the best grid search configurations (range: $\times 0.1$ to $\times 10$), with 50 trials conducted to systematically explore this refined hyperparameter space. The optimal hyperparameter settings identified through Optuna optimization are presented in Table 6.
@@
$\textsc{DIVER}_{\text{Small}/\text{I}/{0.1s}}$ iEEG 0.1s 4.91e-03 3.75e-06
@@ B.8 Finetuning Setup
Unlike Neuroprobe, for which PopT and BrainBERT baselines are provided as benchmarks (2024neuroprobe), the MAYO seizure dataset is an extended dataset that we constructed to match Brant’s minimum input length, and therefore required training on the baseline models.BrainBERT was set at lr = 1e-3 for the classifier with AdamW, as in the original paper (wang2023brainbert), with batch size 32. The features in time [l-5:l+5] were concatenated along the channel dimension. Brant was set at lr = 1e-4 for the classifier and 1e-7 for encoder layers, with betas=(0.9, 0.999), eps=1e-8, batch size 4 and Adam, same as their publicly released code. We could not evaluate PopT on MAYO because LPI coordinates are not available. Brant was pretrained with a 6 s patch and a total of 15 patches (90 s), so we were therefore unable to evaluate it on the neuroprobe (1 s), and there is a mismatch with the original pretrained context in MAYO (6 s).
@@
For all EEG downstream tasks, we use the same optimizer (AdamW) and learning rate scheduler (cosine annealing) as described in the iEEG finetuning configuration. The base learning rate is set to 2.00e-4, weight decay to 3.00e-1, and batch size to 64. We perform full finetuning without employing multi-lr strategies, applying the same learning rate to both the backbone and classifier. The classifier consists of a 3-layer MLP with ELU activation functions, where the first hidden layer has width $T\times 200$ ($T$ is the sample duration in seconds), the second hidden layer has width 200, and the output layer dimension matches the number of classes. Dropout rate is set to 0.1, label smoothing to 0.1, and gradient clipping value to 1.0. All models are trained for 50 epochs without early stopping. For model selection, we use the epoch that achieves the best validation performance (AUROC for binary classification, F1 score for multi-class classification), which is then evaluated on the test set. All results are reported as mean $\pm$ standard deviation across 5 random seeds (41, 42, 43, 44, 45). For the one-to-one comparison experiments with CBraMod (Table 24), we use task-specific hyperparameters to match CBraMod’s training conditions: on Mumtaz2016(Mumtaz2016), linear probing uses learning rate 5.00e-6 and weight decay 6.25e-6, while multi-lr and linear classifier configurations use learning rate 6.25e-5 and weight decay 6.25e-6.
@@ C.1 Scaling Law Experiment Details
Pretraining loss curves for the trained models are presented in Figure 4 for the $\textsc{DIVER}_{\text{-}/\text{I}/1s}$ model family, Figure 5 for the $\textsc{DIVER}_{\text{-}/\text{I}/0.1s}$ model family and Figure 6 for the $\textsc{DIVER}_{\text{-}/\text{E}/1s}$ model family. For the iEEG models, we trained separate instances for each epoch to obtain the corresponding loss curves. In contrast, for EEG models, we trained all models with a fixed maximum of 32 epochs and extracted test loss values at the relevant epoch checkpoints for scaling analysis. This difference in training procedure was necessitated by computational constraints, as EEG model training requires substantially longer wall-clock time. Consequently, the iEEG plots display epoch-specific loss curves(4 and 5), while the EEG plot presents all scaling curves within a single figure(Figure 6). Importantly, this methodological difference affects the learning rate schedule: we employed cosine annealing with warmup, where the decay schedule depends on the total number of training epochs. Therefore, extracting the loss at epoch 2 from a model trained for 32 epochs differs from the loss at epoch 2 of a model trained for only 2 epochs, as the learning rate trajectories diverge under these configurations.
@@
Data-constrained Scaling Law We trained a total of 31 models with varying parameter counts and numbers of training epochs, while keeping the dataset fixed. The hidden dimension was fixed to 12 layers. Since different granularity led to different scaling behavior, we experimented on two granularity conditions and fitted them separately. The empirical isoLoss contours in Figure 2(j) show less smoothness compared to the original scaling law paper(data_constrained_scaling_muennighoff2023scaling), primarily due to sampling density. While the original study used  93 model configurations with dense sampling across all loss ranges, we evaluated  30 configurations. Despite this visual difference, our empirical isoLoss contours (Figure 2(j) and Figure 7, left panel) align well with the predicted contours (Figure 2(j) and Figure 7, right panel), demonstrating that data-constrained scaling laws generalize to neural data.
C.2 Extended Kaplan (kaplan2020scaling) Scaling Law Results for iEEG
@@ C.3 Data-Constrained Scaling Law Fitting Results for iEEG
The exponents $\alpha$ and $\beta$ govern the marginal benefit of scaling parameters and data, respectively. Our values ($\alpha=0.377$, $\beta=0.350$) indicate that increasing model size and adding data yield comparable contributions to overall performance improvements. Compared to prior results in language domain ($\alpha=0.348$, $\beta=0.366$; (hoffmann2022chinchilla)), our fitted exponents show similar values.
@@
C.4 Extended Kaplan (kaplan2020scaling) Scaling Law Results for EEG

### v2

@@ 5 Conclusion
This work presents the first systematic investigation of scaling laws for Ephys foundation models (EFMs), across both scalp-EEG and iEEG modalities. We introduce DIVER-1, a family of EEG/iEEG foundation models ranging from 13M to 1.82B parameters, designed with electrophysiology-specific inductive priors. Our analysis shows that EFMs operate in a data-constrained regime with scaling behavior that fundamentally differs from language models. Through unprecedented scaling across data, model size, and compute, DIVER-1 achieves state-of-the-art performance across various iEEG and EEG tasks.
Our findings underscore that data, not model capacity, is the critical bottleneck for EFM progress. Substantial quantities of ephys data remain untapped in laboratories worldwide. We therefore advocate for open science and community-driven efforts to curate large-scale public datasets. Importantly, our scaling analysis shows that competitive EFMs can be trained even under limited computational budgets by prioritizing longer training durations with small models, lowering the barrier to foundation model development for academic groups with modest GPU resources.
Looking ahead, these findings establish the foundation for a new generation of EFM development by demonstrating that Ephys requires tailored scaling strategies, rather than direct application of scaling laws from other domains. While the precise scaling law parameters and optimal ratios we derived are tied to our particular architecture and pretraining objective, the central insight—that data-constrained regimes favor extended training over increased model size—constitutes a broadly applicable principle for future EFM development.
6 Impact Statement
This work presents controlled scaling experiments across compute, data, and model capacity for electrophysiology foundation models (EFMs) in both scalp EEG and intracranial EEG (iEEG). We identify a data-constrained regime in which compute-efficient performance is achieved by balancing model size with extended multi-epoch training, rather than by increasing parameter count alone. These findings provide practical guidance for allocating limited computational resources when training EFMs.
We plan to release our pretrained models under an open-science framework to support reuse in brain–computer interface and clinical applications. Our conclusions are specific to EEG and iEEG under the modeling and training settings studied here, and scaling behavior may differ for other modalities such as MEG or fMRI.
@@ A.1 Ephys Data Foundation Models
Ephys decoding has progressed from pipelines that coupled hand-crafted features with classical classifiers to end-to-end deep architectures that learn task-relevant representations directly from raw signals. Early EEG studies relied on feature engineering (e.g., band-power, CSP) paired with SVMs or LDA; subsequent work introduced convolutional backbones and sequence models that absorb spectral–temporal patterns with minimal preprocessing. Building on advances in self-supervised learning (SSL) and “foundation” paradigms, recent efforts pretrain large models on heterogeneous, weakly labeled or unlabeled EEG/iEEG corpora and adapt them to diverse downstream tasks (e.g., event detection, cognitive state decoding, BCI control). Representative lines include large-scale EEG transformers and montage-aware encoders (e.g., EEGFormer(Chen et al., 2024), NEURO-GPT(Cui et al., 2024), LaBraM(Jiang et al., 2024)), intracranial representation learners (e.g., BrainBERT(Wang et al., 2023), foundation models for intracranial neural signals), and broader neuro-sequence backbones emphasizing transfer and robustness (e.g., CBraMod(Wang et al., 2024c)).
@@ A.2 Neural Scaling Law
Kaplan (Kaplan et al., 2020) scaling law provides a principled framework for predicting model performance as a function of model size and dataset size for large language models. They demonstrated that language model cross-entropy loss follows smooth power-law relationships with respect to model parameters ($N$) and training data ($D$). Concretely, they proposed a relation of the form
@@
Data-constrained scaling law. In scientific domains, including Ephys, unique high-quality data are inherently limited. To address this, Muennighoff et al. (2023) extended the scaling framework to data-constrained regimes, where models must repeatedly train on the same corpus for multiple epochs. They proposed a modified law
@@
Compared to the Chinchilla law (Hoffmann et al., 2022), which assumes abundant data and one training epoch, this formulation makes epoch count itself a central axis of scaling. This perspective is crucial for domains constrained by limited samples, such as EEG and iEEG, and it guides how compute should be allocated between model size and additional training passes.
Finally, the data-constrained scaling law also connects scaling analysis to more fundamental theory. Recent work has suggested that the exponents governing power-law scaling, such as $\alpha$ and $\beta$, are intimately connected to the intrinsic dimension of the underlying data manifold (Sharma and Kaplan, 2022). In particular, it has been argued that the data exponent $\beta$ can be interpreted in terms of an effective dimension $d$ via an approximate inverse relation of the form:
@@ B.2 Pretraining Dataset
AJILE12 (Peterson et al., 2022) ECoG 12 124,423 1,282 1,000
@@
TUEG (Obeid and Picone, 2016) EEG 10,874 422,036 23,178 250–512
HBN (Shirazi et al., 2024) EEG 2,782 61,703 572 500
NCHSDB (Lee et al., 2022) EEG 3,673 163,146 26,055 256–512
PEERS (Kahana et al., 2023) EEG 364 870,447 6,964 500
@@ B.3 Downstream Task and Dataset Overview
iEEG tasks. We evaluated on 15 tasks from the Neuroprobe (LITE) benchmark (Zahorodnii et al., 2025), including visual perception (frame brightness, optical flow, face detection), auditory processing (volume, pitch, delta volume), and language processing (speech decoding, word prediction, onset detection, part-of-speech tagging). The Neuroprobe dataset contains depth electrode recordings from 6 subjects with 109-120 channels per subject, originally sampled at 2048Hz and was resampled to 500Hz to match our pretraining configuration. There are both binary-label and multi-label task options for Neuroprobe.  [^2] The multi-label setting is currently available only in the released code and has not yet been documented in the paper. In the multi-label configuration, the speech, onset, and head word position tasks remain binary, the part-of-speech task uses 6 labels, and the remaining tasks use 3 labels. Throughout this paper, unless explicitly stated otherwise as multi-label, all reported Neuroprobe results correspond to the binary-label setting. Also, we evalauted on the seizure detection test on MAYO dataset (modified the dataset in kaggle challenge (Bbrinkm and Cukierski, 2014)). The original dataset consists of 1 s samples, but to match the minimum patch length of Brant, one of the baseline models, we concatenated samples in temporal order to create 6 s samples. We evaluated the model separately for each of the 8 participants. To address the issue of having more test data than training data, we swapped the train and test sets for each participant.
EEG tasks. We evaluated on three EEG benchmarks: FACED (emotion recognition from 32-channel EEG, 9-class classification, 123 subjects)(Chen et al., 2023), PhysioNet-MI (motor imagery with 64 channels, 4-class classification, 109 subjects)(Goldberger et al., 2000), and MentalArithmetic (mental stress detection with 20 channels reduced to 19, 2-class classification, 36 subjects)(Zyma et al., 2019). Sampling rates were standardized to 500Hz across all datasets to ensure consistency with our pretraining setup.
@@ B.4 Pretraining Setup and Model Scaling
We varied the model size by modifying the hidden dimension of the transformer, resulting in sizes of 13M, 51M, 115M, 203M, 813M, 1.83B parameters, while keeping the depth fixed at 12 layers. This capacity adjustment leverages the benefits of $\mu$ parameterization for stable training across different model sizes. DIVER-1 was implemented on the Python 3.12.3 and Pytorch 2.6.0 + cuda version 12.4. To enhance training efficiency, we employed DeepSpeed ZeRO Stage 2, BF16 precision. Optimization was performed using a custom implementation of the DeepSpeed’s MuAdam optimizer (Yang et al., 2022) with utilizing DeepSpeed’s FusedAdam backend (Rasley et al., 2020) for computational efficiency and learning rate calibration. A cosine annealing learning rate scheduler with warm-up restarts was applied, with cycle length matching the total training steps and minimum learning rate set to 0.01× the initial rate.
@@ B.5 DIVER Architecture and Pretext Task Hyperparamter Setting
Positional Embedding Channel type dimension $d_{\text{model}}/4$
@@
Stage2: Optuna Optimization We further refined the hyperparameters using Optuna(Akiba et al., 2019) or bayesian hyperparameter optimization. The search space was defined as $\pm 1$ order of magnitude around the best grid search configurations (range: $\times 0.1$ to $\times 10$), with 50 trials conducted to systematically explore this refined hyperparameter space. The optimal hyperparameter settings identified through Optuna optimization are presented in Table 7.
@@
$\textsc{DIVER}_{\text{Small}/\text{I}/{\color[rgb]{0,0,0}0.1s}}$ iEEG 0.1s 4.91e-03 3.75e-06
@@ B.8 Finetuning Setup
Unlike Neuroprobe, for which PopT and BrainBERT baselines are provided as benchmarks (Zahorodnii et al., 2025), the MAYO seizure dataset is an extended dataset that we constructed to match Brant’s minimum input length, and therefore required training on the baseline models.BrainBERT was set at lr = 1e-3 for the classifier with AdamW, as in the original paper (Wang et al., 2023), with batch size 32. The features in time [l-5:l+5] were concatenated along the channel dimension. Brant was set at lr = 1e-4 for the classifier and 1e-7 for encoder layers, with betas=(0.9, 0.999), eps=1e-8, batch size 4 and Adam, same as their publicly released code. We could not evaluate PopT on MAYO because LPI coordinates are not available. Brant was pretrained with a 6 s patch and a total of 15 patches (90 s), so we were therefore unable to evaluate it on the neuroprobe (1 s), and there is a mismatch with the original pretrained context in MAYO (6 s).
@@
For all EEG downstream tasks, we use the same optimizer (AdamW) and learning rate scheduler (cosine annealing) as described in the iEEG finetuning configuration. The base learning rate is set to 2.00e-4, weight decay to 3.00e-1, and batch size to 64. We perform full finetuning without employing multi-lr strategies, applying the same learning rate to both the backbone and classifier. The classifier consists of a 3-layer MLP with ELU activation functions, where the first hidden layer has width $T\times 200$ ($T$ is the sample duration in seconds), the second hidden layer has width 200, and the output layer dimension matches the number of classes. Dropout rate is set to 0.1, label smoothing to 0.1, and gradient clipping value to 1.0. All models are trained for 50 epochs without early stopping. For model selection, we use the epoch that achieves the best validation performance (AUROC for binary classification, F1 score for multi-class classification), which is then evaluated on the test set. All results are reported as mean $\pm$ standard deviation across 5 random seeds (41, 42, 43, 44, 45). For the one-to-one comparison experiments with CBraMod (Table 25), we use task-specific hyperparameters to match CBraMod’s training conditions: on Mumtaz2016(Mumtaz, 2016), linear probing uses learning rate 5.00e-6 and weight decay 6.25e-6, while multi-lr and linear classifier configurations use learning rate 6.25e-5 and weight decay 6.25e-6.
@@ C.1 Scaling Law Experiment Details
Pretraining loss curves for the trained models are presented in Figure 5 for the $\textsc{DIVER}_{\text{-}/\text{I}/1s}$ model family, Figure 6 for the $\textsc{DIVER}_{\text{-}/\text{I}/0.1s}$ model family and Figure 7 for the $\textsc{DIVER}_{\text{-}/\text{E}/1s}$ model family. For the iEEG models, we trained separate instances for each epoch to obtain the corresponding loss curves. In contrast, for EEG models, we trained all models with a fixed maximum of 32 epochs and extracted test loss values at the relevant epoch checkpoints for scaling analysis. This difference in training procedure was necessitated by computational constraints, as EEG model training requires substantially longer wall-clock time. Consequently, the iEEG plots display epoch-specific loss curves(5 and 6), while the EEG plot presents all scaling curves within a single figure(Figure 7). Importantly, this methodological difference affects the learning rate schedule: we employed cosine annealing with warmup, where the decay schedule depends on the total number of training epochs. Therefore, extracting the loss at epoch 2 from a model trained for 32 epochs differs from the loss at epoch 2 of a model trained for only 2 epochs, as the learning rate trajectories diverge under these configurations.
@@
Data-constrained Scaling Law We trained a total of 31 models with varying parameter counts and numbers of training epochs, while keeping the dataset fixed. The hidden dimension was fixed to 12 layers. Since different granularity led to different scaling behavior, we experimented on two granularity conditions and fitted them separately. The empirical isoLoss contours in Figure 3(j) show less smoothness compared to the original scaling law paper(Muennighoff et al., 2023), primarily due to sampling density. While the original study used  93 model configurations with dense sampling across all loss ranges, we evaluated  30 configurations. Despite this visual difference, our empirical isoLoss contours (Figure 3(j) and Figure 8, left panel) align well with the predicted contours (Figure 3(j) and Figure 8, right panel), demonstrating that data-constrained scaling laws generalize to neural data.
C.2 Extended Kaplan (Kaplan et al., 2020) Scaling Law Results for iEEG
@@ C.3 Data-Constrained Scaling Law Fitting Results for iEEG
The exponents $\alpha$ and $\beta$ govern the marginal benefit of scaling parameters and data, respectively. Our values ($\alpha=0.377$, $\beta=0.350$) indicate that increasing model size and adding data yield comparable contributions to overall performance improvements. Compared to prior results in language domain ($\alpha=0.348$, $\beta=0.366$; (Hoffmann et al., 2022)), our fitted exponents show similar values.
@@
C.4 Extended Kaplan (Kaplan et al., 2020) Scaling Law Results for EEG